# The Structure of Hypergraphs Arising in Cellular Mobile Communication Systems

### v3

@@ 7 Nonrealizability of the Hypergraph $K_{1,5}$ When $\alpha=3$
If the path loss exponent $\alpha=3$, then the hypergraph $K_{1,4}$ is realizable.
Proof: Consider the wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ where $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$, $\alpha=3$, $\beta=1$, $s_{0}$ is placed at the origin, and the remaining stations $s_{i}(i=1,\ldots,4)$ are placed uniformly on the unit circle in clockwise order. Then, $\{s_{0},s_{i}\}$ is a minimal forbidden set for all $i=1,\ldots,4$. The energy at one of the stations in $W=\{s_{1},s_{2},s_{3},s_{4}\}$, say $s_{1}$, due to transmissions by the remaining stations in $W$, is given by
@@
If the path loss exponent $\alpha=3$, then the hypergraph $K_{1,5}$ is not realizable.
Proof: The proof consists of two distinct segments. In the first segment, a proof by contradiction assumes the realizability of the star $K_{1,5}$. A function $f$ is introduced to represent the interference at one of the stations due to transmissions by two of its close neighbors, establishing that its value is less than $1$ at some point in its domain. Transitioning to the second segment, specifically Lemma 18, it is then demonstrated that the function $f$ fails to adhere to the aforementioned property.
By way of contradiction, suppose there exists a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ that generates the hypergraph $(S,\mathcal{E})=K_{1,5}$, where $S=\{s_{0},s_{1},\ldots,s_{5}\}$ and $\mathcal{E}$ is the collection $\{\{s_{0},s_{i}\}:i\in[5]\}$ of five hyperedges. Recall that $\alpha=3$. By Lemma 12, it can be assumed that $\beta=1$. Since a translation of the Euclidean plane is a rigid motion and preserves distances, it can be assumed that the station $s_{0}$ is located at the origin.
$E(W-\{w\},w)=\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{3}}<1.$
We will arrive at a contradiction by proving that this inequality is violated for every set $W$ of five locations in the unit ball (minus the origin).
The points in $W$ can be pushed radially outward to the boundary of the unit ball, i.e. to the unit circle, and by Lemma 11 this transformation does not decrease the distance between any two points in $W$ and hence preserves feasibility of $W$. Thus, we can assume that the hypergraph $K_{1,5}$ is generated by a wireless network $\langle S,\alpha,\beta\rangle$, such that $S=\{s_{0},s_{1},\ldots,s_{5}\}$, $\alpha=3$, $\beta=1$, and the set of points $W=\{s_{1},\ldots,s_{5}\}$ lies on the unit circle.
Figure 1: Five stations on the unit circle, with stations $s_{1},s_{2}$ and $s_{5}$ close enough to each other that $\delta_{1}+\delta_{5}=2\pi/15$. Relevant to proof of Theorem 17.
$f(\delta_{1})=\frac{1}{d(s_{1},s_{2})^{3}}+\frac{1}{d(s_{1},s_{5})^{3}},$
where the domain for $f$ is $[0,2\pi/15]$. By Proposition 18, the minimum value of $f(\delta_{1})$ is greater than $1$. Because the reception threshold $\beta$ is equal to $1$, the subset $\{s_{1},s_{2},s_{5}\}$ is a forbidden subset. Hence, its superset $W$ is also a forbidden set, contradicting the fact that $W$ is an independent set.
@@
Let $f(\delta_{1})$ be the function
$f(\delta_{1}):=\frac{1}{d(s_{1},s_{2})^{3}}+\frac{1}{d(s_{1},s_{5})^{3}},$
Define the function $g(\delta)=d(s_{1},s_{2})^{-3}+d(s_{1},s_{5})^{-3}$, where $s_{1}=(1,0)$, $s_{2}=(1,2\pi/5+\delta)$, $s_{5}=(1,2\pi/5-\delta)$, and the domain is $[-\pi/15,\pi/15]$. Because $f(\delta_{1})=g(\delta+\pi/15)$, it suffices to show that the minimum value of $g(\delta)$ is greater than $1$. To prove that $g(\delta)>1$ over its domain, it suffices to prove that $g$ is convex and $g^{\prime}(0)=0$, implying $g$ has a global minimum at $\delta=0$, and furthermore, that $g(0)>1$.
Let $g_{1}(\delta)$ and $g_{2}(\delta)$ denote the first term $d(s_{1},s_{2})^{-3}$ and the second term $d(s_{1},s_{5})^{-3}$, respectively, in the expression for $g(\delta)$, so that $g(\delta)=g_{1}(\delta)+g_{2}(\delta)$. It can be verified that
$g_{1}(\delta)=\frac{1}{\sqrt{8}}\left(1-\cos\left(\frac{2\pi}{5}+\delta\right)\right)^{-3/2}$
$g_{1}^{\prime}(\delta)=\frac{-3}{4\sqrt{2}}\sin\left(\frac{2\pi}{5}+\delta\right)\left[1-\cos\left(\frac{2\pi}{5}+\delta\right)\right]^{-5/2}$
$g_{1}^{\prime\prime}(\delta)=\frac{-3}{4\sqrt{2}}h_{1}(\delta)h_{2}(\delta),$
where the functions $h_{1}$ and $h_{2}$ are defined to be
$h_{1}(\delta)=\left[1-\cos\left(\frac{2\pi}{5}+\delta\right)\right]^{-7/2}$
$h_{2}(\delta)=\cos\left(\frac{2\pi}{5}+\delta\right)-1-\frac{3}{2}\sin^{2}\left(\frac{2\pi}{5}+\delta\right).$
For all $\delta$ in domain, $\cos\left(\frac{2\pi}{5}+\delta\right)-1<0$ and $\sin^{2}\left(\frac{2\pi}{5}+\delta\right)>0$, whence $h_{2}(\delta)<0$. Thus, $g_{1}^{\prime\prime}(\delta)>0$ for all $\delta$. A similar proof shows that $g_{2}^{\prime\prime}(\delta)>0$ for all $\delta$. Hence, $g(\delta)=g_{1}(\delta)+g_{2}(\delta)$ is convex over its domain.
Furthermore, it can be verified that $g_{1}^{\prime}(0)+g_{2}^{\prime}(0)=0$, whence $g(0)=0$. Finally, it can be shown that $g(0)=\frac{2}{\sqrt{25-10\sqrt{5}}}$, which is greater than $1$. This completes the proof.
8 Nonrealizability of the Hypergraph $K_{1,4}$ When $\alpha=2$
In the present section, we illustrate that the proof techniques given in the previous section can be utilized to obtain results for other values of the path loss exponent $\alpha$. More specifically, we show that when $\alpha=2$, the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=3$. Compared to the previous section, the proof here is brief because the ingredients in the proof are the same as before, mutatis mutandis. In the rest of this section, assume $\alpha=2$.
@@ 8 Nonrealizability of the Hypergraph $K_{1,4}$ When $\alpha=2$
If the path loss exponent $\alpha=2$, then the hypergraph $K_{1,3}$ is realizable.
Proof: Let $S=\{s_{0},s_{1},s_{2},s_{3}\}$ be a set of stations such that $s_{0}=(0,0)$, and $s_{i}$ has polar coordinates $(r,\theta_{i})=(1,2\pi i/3)$ $(i=1,2,3)$. Thus, the latter three points are spaced uniformly on the unit circle. Observe that the distance between $s_{0}$ and $s_{i}$ is at most $1$, for all $i$. Let $\beta=1$ and recall $\alpha=2$. It follows that hypergraph generated by the wireless network $\langle S,\alpha,\beta\rangle$ is $(S,\mathcal{E})$, where $\mathcal{E}$ contains three edges $\{s_{0},s_{1}\}$, $\{s_{0},s_{2}\}$, and $\{s_{0},s_{3}\}$.
Let $W=\{s_{1},s_{2},s_{3}\}$. Pick any $w\in W$. It can be verified that the energy $E(W-\{w\},w\}$ at $w$ due to transmissions by the other two stations is $2/3$, which is less than $1$. Hence, $W$ is feasible, which implies that there are no other minimal forbidden sets other than the three edges mentioned above. Thus, the hypergraph generated by the wireless network $\langle S,\alpha,\ \beta\rangle$ is $K_{1,3}$.
@@
If the path loss exponent $\alpha=2$, then the hypergraph $K_{1,4}$ is not realizable.
Proof: By way of contradiction, suppose there exists a set of locations $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$ in the Euclidean plane, with $\alpha=2$ and $\beta=1$ such that the wireless network $\langle S,\alpha,\beta\rangle$ generates the hypergraph $K_{1,4}$. Without loss of generality, we may assume $s_{0}$ is the center vertex of the star $K_{1,4}$ and we may assume $s_{0}$ is at the origin. Thus, $W=\{s_{1},s_{2},s_{3},s_{4}\}$ is feasible and the points in $W$ lie on the unit ball (minus the origin). Also, the distance between any two distinct points in $W$ is greater than $1$. By Lemma 11, the points in $W$ can be pushed radially outward to the boundary of the unit circle, and since this only increases the distance between points in $W$, the resulting set of four points on the unit circle is a feasible set.
Because rotations preserve distances, it can be assumed without loss of generality that the polar coordinate of $s_{1}$ is $(1,0)$. Also, the remaining three points can be assumed to be on the unit circle in counterclockwise order, with $\theta_{i}$ denoting the angle between $s_{i}$ and $s_{i+1}$; see Figure 2. Thus, $s_{2}$ has polar coordinates $(1,\theta_{1})$, $s_{3}$ has polar coordinates $(1,\theta_{1}+\theta_{2})$ and $s_{4}$ has polar coordinates $(1,\theta_{1}+\theta_{2}+\theta_{3})$, which we shall also denote by $(1,-\theta_{4})$.
$\theta_{1}$$\theta_{4}$$x$$y$$s_{2}$$s_{1}$$\theta_{1}=\frac{\pi}{3}+\delta_{1}$$\theta_{4}=\frac{\pi}{3}+\delta_{4}$$s_{3}$$s_{4}$
Figure 2: Four stations on the unit circle, with stations $s_{1},s_{2}$ and $s_{4}$ close enough to each other that $\delta_{1}+\delta_{4}=\pi/3$. Relevant to proof of Theorem 20.
The distance between any two points in $W$ is greater than $1$, whence $\theta_{i}>\pi/3$ for all $i$. It follows that the sum of any three $\theta_{i}$’s is more than $\pi$, whence each $\theta_{i}$ is less than $\pi$. Thus, we have that $\pi/3<\theta_{i}<\pi$, for all $i$.
Define $\delta_{i}=\theta_{i}-\pi/3$ $(i=1,\ldots,4)$. Then, $30<\delta_{i}<2\pi/$ 3, for all $i$. The sum of all four $\delta_{i}$’s is $32\pi/$ 3. It follows that there exists some two consecutive $\delta_{i}$’s whose sum is at most $\pi/3$. Without loss of generality, suppose these two consecutive $\delta_{i}$’s are $\delta_{1}$ and $\delta_{4}$. So, $\delta_{1}+\delta_{4}\leq\pi/3$. It will be proved that the energy at $s_{1}$ due to transmissions by $\{s_{2},s_{4}\}$ is at least $1$ even if $s_{2}$ and $s_{4}$ are as far apart as possible from $s_{1}$, i.e. even if $\delta_{1}+\delta_{4}=\pi/3$.
Define
$f(\delta_{1})=\frac{1}{d(s_{1},s_{2})^{2}}+\frac{1}{d(s_{1},s_{4})^{2}},$
where the domain for $f$ is $[0,\pi/3]$, $s_{1}$ has polar coordinates $(1,0)$, $s_{2}$ is at $(1,\pi/3+\delta_{1})$, $s_{4}$ is at $(1,-\pi/3-\delta_{4})$, and $\delta_{1}+\delta_{4}=\pi/3$. We may assume $s_{4}$ is at $(1,\pi/3+\delta_{4})$ since the distance to $s_{1}$ is the same if its angle is negated. Thus, we have function $f$ of a single variable $\delta_{1}$ and it suffices to show that $f(\delta_{1})\geq 1$ for all $\delta_{1}$ in the domain. For this would imply that the energy $E(\{s_{2},s_{4}\},s_{1})$ at $s_{1}$ due to transmissions by $s_{2}$ and $s_{4}$ is at least $1$, and hence for any placement of $s_{3}$ the energy $E(\{s_{2},s_{3},s_{4}\},s_{1})>1$. This would prove that $W=\{s_{1},s_{2},s_{3},s_{4}\}$ is a forbidden set, contradicting feasibility of $W$.
Let $\delta=\delta_{1}-\pi/6$ and define $g(\delta):=f(\delta_{1})$. Then
$g(\delta)=\frac{1}{2(1+\sin\delta)}+\frac{1}{2(1-\sin\delta)},$
where the domain of $g$ is $[-\pi/6,\pi/6]$. The minimum value of $f$ over its domain is equal to the minimum value of $g$ over its domain. Observe that $g$ can be written as the sum of two functions $g=g_{1}+g_{2}$, where $g_{1}(\delta)=\frac{1}{2(1+\sin\delta)}$. It can be shown that
$g_{1}^{\prime}(\delta)=\frac{-\cos\delta}{2(1+\sin\delta)^{2}},$
and the second derivative
$g_{1}^{\prime\prime}(\delta)=\frac{1}{2(1+\sin\delta)^{4}}\left[1+\sin\delta\right]\left[(1+\sin\delta)(\sin\delta)+2\cos^{2}\delta\right]$
is a product of three terms. It is clear that the first and second terms are each positive over the domain. The third term is positive if and only if $\sin^{2}\delta-\sin\delta<2$, which is true over the domain. Thus, $g_{1}^{\prime\prime}$ is positive over the domain. Similarly, it can be shown that $g_{2}^{\prime\prime}$ is positive over the domain. It follows that their sum $g^{\prime\prime}$ is positive over the domain, which implies that $g$ is convex. Also, $g^{\prime}(0)=0$. Hence, $g$ attains its minimum value at $\delta=0$. Since $g(0)=1$, the assertion follows.
9 Nonrealizability of the Hypergraph $K_{1,3}$ When $\alpha=1$
In this section, we consider the case where the path loss exponent $\alpha=1$. Motivated by results on the performance guarantees of distributed algorithms in wireless networks, we determine the maximal value of $r$ such that the $2$-uniform hypergraph $K_{1,r}$ is realizable. This value is shown to be equal to $2$.
Lemma 21.
If the path loss exponent $\alpha=1$, then the hypergraph $K_{1,2}$ is realizable.
Proof: Consider the wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$, where $S=\{s_{0},s_{1},s_{2}\}$, $\alpha=1$, and $\beta=1$, and in rectangular form the locations of the stations are $s_{0}=(0,0)$, $s_{1}=(1,0)$, and $s_{2}=(-1,0)$. Let $H=(S,\mathcal{E})$ be the hypergraph generated by the wireless network $\mathcal{A}$. Because the distance $d(s_{0},s_{i})\leq 1$, $\{s_{0},s_{i}\}\in\mathcal{E}$ ($i=1,2$). The energy at $s_{1}$ due to transmission by $s_{2}$ is given by
$E(\{s_{2}\},s_{1})=\frac{1}{d(s_{1},s_{2})^{\alpha}}=\frac{1}{2^{1}}<1.$
The proof of the next result on the nonrealizability of $K_{1,3}$ consists of two segments. The first segment is similar to the proof of the initial segment above and hence will be brief. The second segment is given in Lemma 23 below.
Theorem 22.
If the path loss exponent $\alpha=1$, then the hypergraph $K_{1,3}$ is not realizable.
@@ 9 Nonrealizability of the Hypergraph $K_{1,3}$ When $\alpha=1$
Proof: By way of contradiction, suppose the hypergraph $K_{1,3}$ is realizable. Then, using the same arguments as the proof of Theorem 17, it can be shown that there exist three locations $s_{1}$, $s_{2}$, and $s_{3}$ on the unit circle, with polar coordinates $s_{1}=(1,0)$, $s_{2}=(1,\theta_{1})$, and $s_{3}=(1,-\theta_{3})$, where $\frac{\pi}{3}<\theta_{i}<\frac{4\pi}{3}$ for all $i$, satisfying the following two conditions: (i) $\theta_{1}+\theta_{3}=\frac{4\pi}{3}$, and (ii) the interference at $s_{1}$ due to transmissions by its two neighbors $s_{2}$ and $s_{3}$ is tolerable in the sense that
$\frac{1}{d(s_{1},s_{2})}+\frac{1}{d(s_{1},s_{3})}<1.$
This inequality holds also if the angle of $s_{3}$ is negated, so let $s_{3}=(1,-\theta_{3})$. Define $\delta=\theta_{1}-\frac{2\pi}{3}$. Then, $s_{1}=(1,0)$, $s_{2}=(1,\delta+\frac{2\pi}{3})$ and $s_{3}=(1,\delta-\frac{2\pi}{3})$. Define
$g(\delta)=\frac{1}{d(s_{1},s_{2})}+\frac{1}{d(s_{1},s_{3})}$
with domain $\delta\in[\frac{-\pi}{3},\frac{\pi}{3}]$. It follows that $g(\delta)<1$ for some $\delta$ in the domain. This is a contradiction because, by Lemma 23, $g(\delta)>1$ for all $\delta$ in the domain.
Lemma 23.
Let $g(\delta)$ be the function
$g(\delta)=\frac{1}{d(s_{1},s_{2})}+\frac{1}{d(s_{1},s_{3})}$
where the points $s_{1}$, $s_{2}$, and $s_{3}$ have polar coordinates $s_{1}=(1,0)$, $s_{2}=(1,\delta+\frac{2\pi}{3})$, and $s_{3}=(1,\delta-\frac{2\pi}{3})$, and the domain is $-\frac{\pi}{3}\leq\delta\leq\frac{\pi}{3}$. Then, the minimum value of $g(\delta)$ is greater than $1$.
Proof: Computing the distances $d(s_{1},s_{2})$ and $d(s_{1},s_{3})$ using the Pythagorean theorem, one obtains that
$g(\delta)=\frac{1}{\sqrt{2+\cos\delta+\sqrt{3}\sin\delta}}+\frac{1}{\sqrt{2+\cos\delta-\sqrt{3}\sin\delta}}.$
Denote the first and second terms of the right hand side by $g_{1}(\delta)$ and $g_{2}(\delta)$, respectively, so that $g(\delta)=g_{1}(\delta)+g_{2}(\delta)$. It can be verified that $g^{\prime}(0)=g_{1}^{\prime}(0)+g_{2}^{\prime}(0)=-\frac{1}{6}+\frac{1}{6}=0$, whence $\delta=0$ is a critical point. Also, $g(0)=\frac{2}{\sqrt{3}}>1$. To complete the proof, it suffices to show that $g_{1}^{\prime\prime}(\delta)>0$ and $g_{2}^{\prime\prime}(\delta)>0$ for all $\delta$, whence $g$ is convex.
By differentiating $g_{1}$, one obtains that
$g_{1}^{\prime}(\delta)=\frac{\sin\delta-\sqrt{3}\cos\delta}{2}\cdot\frac{1}{\left[t_{1}(\delta)\right]^{3/2}},$
where $t_{1}(\delta):=2+\cos\delta+\sqrt{3}\sin\delta$. By differentiating again, one obtains that
$g_{1}^{\prime\prime}(\delta)=\frac{1}{2\left[t_{1}(\delta)\right]^{5/2}}\cdot\left\{\frac{9}{2}+t_{2}(\delta)+t_{3}(\delta)\right\},$
where $t_{2}(\delta):=\cos^{2}\delta+2\cos\delta$, and $t_{3}(\delta):=2\sqrt{3}\sin\delta-\sqrt{3}\cos\delta\sin\delta$. It will be shown that the numerator and denominator of $g_{1}^{\prime\prime}(\delta)$ are both positive for all $\delta$. Because $\cos\delta>0$ for all $\delta$ (in the domain), $t_{2}(\delta)>0$ for all $\delta$. Also, $|t_{3}(\delta)|\leq\sqrt{3}\cdot\frac{\sqrt{3}}{2}\cdot 2=3$. Thus, the numerator of $g_{1}^{\prime\prime}(\delta)$ is positive. Regarding the denominator of $g_{1}^{\prime\prime}(\delta)$, we now show that $t_{1}(\delta)>0$ for all $\delta$. Observe that $t_{1}^{\prime}(\delta)=\sin\delta+\sqrt{3}\cos\delta$. Simplifying, we obtain that $t_{1}^{\prime}(\delta)\geq 0$ iff $\tan\delta\leq\sqrt{3}$, which is true for all $\delta$ (in the domain). Since $t_{1}(-\frac{\pi}{3})=1>0$ and $t_{1}^{\prime}(\delta)\geq 0$ for all $\delta$, it follows that $t_{1}(\delta)>0$ for all $\delta\in[-\frac{\pi}{3},\frac{\pi}{3}]$. Because the denominator of $g_{1}^{\prime\prime}(\delta)$ is also positive, $g_{1}^{\prime\prime}(\delta)>0$ for all $\delta$. The same proof, mutatus mutandis, gives that $g_{2}^{\prime\prime}(\delta)>0$ for all $\delta$. Hence, $g$ is convex. This proves the assertion.
10 Extensions to Non-integral Values of the Path Loss Exponent $\alpha$

### v4

@@ 7 Nonrealizability of the Hypergraph $K_{1,5}$ When $\gamma=3$
If the path loss exponent $\gamma=3$, then the hypergraph $K_{1,4}$ is realizable.
Proof: Consider the wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ where $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$, $\gamma=3$, $\beta=1$, $s_{0}$ is placed at the origin, and the remaining stations $s_{i}(i=1,\ldots,4)$ are placed uniformly on the unit circle in clockwise order. Then, $\{s_{0},s_{i}\}$ is a minimal forbidden set for all $i=1,\ldots,4$. The energy at one of the stations in $W=\{s_{1},s_{2},s_{3},s_{4}\}$, say $s_{1}$, due to transmissions by the remaining stations in $W$, is given by
@@
Intuitively, as the path loss exponent $\gamma$ increases, the signal attenuates more rapidly with distance, leading to reduced interference at the receiver. This makes the realization of $K_{1,5}$ more likely. Conversely, as $\gamma$ decreases, the likelihood of realizing $K_{1,5}$ diminishes. This observation is captured in Proposition 28. Since $K_{1,4}$ is realizable even at lower values of $\gamma$, such as $\gamma=3$, it remains realizable at higher values, such as $\gamma=4$. In other words, Lemma 19 implies Lemma 15. Similarly, because $K_{1,5}$ is not realizable even at larger values of $\gamma$, such as $\gamma=4$, it is also not realizable at $\gamma=3$:
@@
If the path loss exponent $\gamma\leq 3$, then the hypergraph $K_{1,5}$ is not realizable.
Proof: The assertion follows from Theorem 16 and Proposition 28.
By Theorem 20, when the path loss exponent $\gamma=3$, the hypergraph $K_{1,5}$ is not realizable. Combined with Lemma 19, this leads to the conclusion that for $\gamma=3$, the maximal value of $r$ for which the hypergraph $K_{1,r}$ is realizable is $r=4$.
8 Nonrealizability of the Hypergraph $K_{1,4}$ When $\gamma\leq 2$
In the present section, we show that when $\gamma=2$, the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=3$.
@@ 8 Nonrealizability of the Hypergraph $K_{1,4}$ When $\gamma\leq 2$
If the path loss exponent $\gamma=2$, then the hypergraph $K_{1,3}$ is realizable.
Proof: Let $S=\{s_{0},s_{1},s_{2},s_{3}\}$ be a set of stations such that $s_{0}=(0,0)$, and $s_{i}$ has polar coordinates $(r,\theta_{i})=(1,2\pi i/3)$ $(i=1,2,3)$. Thus, the latter three points are spaced uniformly on the unit circle. Observe that the distance between $s_{0}$ and $s_{i}$ is at most $1$, for all $i$. Let $\beta=1$ and recall $\gamma=2$. It follows that hypergraph generated by the wireless network $\langle S,\gamma,\beta\rangle$ is $(S,\mathcal{E})$, where $\mathcal{E}$ contains three edges $\{s_{0},s_{1}\}$, $\{s_{0},s_{2}\}$, and $\{s_{0},s_{3}\}$.
Let $W=\{s_{1},s_{2},s_{3}\}$. Pick any $w\in W$. It can be verified that the energy $E(W-\{w\},w\}$ at $w$ due to transmissions by the other two stations is $2/3$, which is less than $1$. Hence, $W$ is feasible, which implies that there are no other minimal forbidden sets other than the three edges mentioned above. Thus, the hypergraph generated by the wireless network $\langle S,\gamma,\ \beta\rangle$ is $K_{1,3}$.
Theorem 22.
If the path loss exponent $\gamma\leq 2$, then the hypergraph $K_{1,4}$ is not realizable.
Proof: Fix $\gamma\leq 2$. By way of contradiction, suppose there exists a set of locations $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$ in the Euclidean plane, with $\beta=1$, such that the wireless network $\langle S,\gamma,\beta\rangle$ generates the hypergraph $K_{1,4}$. Without loss of generality, we may assume $s_{0}$ is the center vertex of the star $K_{1,4}$ and we may assume $s_{0}$ is at the origin. Thus, $W=\{s_{1},s_{2},s_{3},s_{4}\}$ is feasible and the points in $W$ lie on the unit ball (minus the origin). Also, the distance between any two distinct points in $W$ is greater than $1$. By Lemma 13, the points in $W$ can be pushed radially outward to the boundary of the unit ball (namely the unit circle), and since this only increases the distance between points in $W$, the resulting set of four points on the unit circle is a feasible set.
Without loss of generality, choose a station in $W$, say $s_{1}$. Because the interference at $s_{1}$ due to senders in $W-\{s_{1}\}$ is tolerable, we have that
$\sum_{s_{j}\in W-\{s_{1}\}}\frac{1}{d(s_{j},s_{1})^{\gamma}}<1.$
The distance between any two stations on the unit circle is at most $2$, so each of the three terms in the sum above is at least $\frac{1}{4}$. By the inequality above, each term in the sum above is less than $\frac{1}{2}$. In particular, if $s_{j}$ is the station in $W-\{s_{1}\}$ that is closest to $s_{1}$, then $\frac{1}{d(s_{j},s_{1})^{\gamma}}<\frac{1}{2}$, i.e. $d(s_{j},s_{1})>\sqrt{2}$. This implies the angle between $s_{1}$ and its nearest neighbor $s_{j}$ is more than $90$ degrees. Since $s_{1}$ was chosen without loss of generality, this conclusion holds for all stations in $W$. In other words, there are four stations on the unit circle such that the angle between any two consecutive stations is more than $90$ degrees, an impossibility.
9 Nonrealizability of the Hypergraph $K_{1,3}$ When $\gamma\leq 1$
In this section, we consider the case where the path loss exponent $\gamma\leq 1$. Motivated by results on the performance guarantees of distributed algorithms in wireless networks, we determine the maximal value of $r$ such that the $2$-uniform hypergraph $K_{1,r}$ is realizable. This value is shown to be equal to $2$.
@@ 9 Nonrealizability of the Hypergraph $K_{1,3}$ When $\gamma\leq 1$
If the path loss exponent $\gamma>0$, then the hypergraph $K_{1,2}$ is realizable.
Proof: Fix $\gamma>0$. Consider the wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$, where $S=\{s_{0},s_{1},s_{2}\}$, $\beta=1$, and in rectangular form the locations of the stations are $s_{0}=(0,0)$, $s_{1}=(1,0)$, and $s_{2}=(-1,0)$. Let $H=(S,\mathcal{E})$ be the hypergraph generated by the wireless network $\mathcal{A}$. Because the distance $d(s_{0},s_{i})\leq 1$, $\{s_{0},s_{i}\}\in\mathcal{E}$ ($i=1,2$). The energy at $s_{1}$ due to transmission by $s_{2}$ is given by
$E(\{s_{2}\},s_{1})=\frac{1}{d(s_{1},s_{2})^{\gamma}}=\frac{1}{2^{\gamma}}<1.$
@@
If the path loss exponent $\gamma\leq 1$, then the hypergraph $K_{1,3}$ is not realizable.
@@
Proof: Let $\gamma\leq 1$. By way of contradiction, suppose the hypergraph $K_{1,3}$ is realizable. Then, using the same arguments as in the proof of Theorem 20, it can be shown that there exists a set $W=\{s_{1},s_{2},s_{3}\}$ of three stations on the unit circle such that $W$ is feasible. The distance between any two points in $W$ is at most $2$. Thus, $d(s_{1},s_{2})\leq 2$ and $d(s_{1},s_{3})\leq 2$. The interference at $s_{1}$ due to transmissions by $s_{2}$ and $s_{3}$ is
$E(\{s_{2},s_{3}\},s_{1})=\frac{1}{d(s_{1},s_{2})^{\gamma}}+\frac{1}{d(s_{1},s_{3})^{\gamma}}\geq\frac{1}{2}+\frac{1}{2}=1.$
This implies $W$ is a forbidden set, a contradiction.
10 Extensions to Non-integral Values of the Path Loss Exponent $\gamma$